OpenStudy (clarence):

My answer for this question was 4*Pi*(x^2+2)... Find the area of the surface obtained by rotating the curve about the y-axis. y=2x^2+4, 0

jimthompson5910 (jim_thompson5910):

You should get a single number and not a variable expression

OpenStudy (clarence):

I should?

jimthompson5910 (jim_thompson5910):

yes, there shouldn't be an x in the final answer

OpenStudy (clarence):

Oh right... I'll check my working out again and will be right back..

OpenStudy (clarence):

\[\frac{ 1 }{ 128 }\pi(644 \sqrt{17}+127arcsinh(4))\] That was what I got, was that what you obtained as well?

jimthompson5910 (jim_thompson5910):

did you integrate this? \[\Large \int_{0}^{1}2\pi x \sqrt{1+(4x)^2}dx\]

OpenStudy (clarence):

Not quite, I integrated this: \[\int\limits_{0}^{1}2\pi(4+2x ^{2})\sqrt{1+16x^2}dx\]

jimthompson5910 (jim_thompson5910):

Here's how I did it \[\Large \int_{0}^{1}2\pi x \sqrt{1+(4x)^2}dx\] \[\Large \int_{0}^{1}2\pi x \sqrt{1+16x^2}dx\] Let u = 1+16x^2 du/dx = 32x ---> x*dx = du/32 if x = 0, then u = 1+16x^2 = 1+16(0)^2 = 1 if x = 1, then u = 1+16x^2 = 1+16(1)^2 = 17 \[\Large \int_{0}^{1}2\pi x \sqrt{1+16x^2}dx\] \[\Large \int_{0}^{1}2\pi\sqrt{1+16x^2}*x*dx\] \[\Large 2\pi\int_{1}^{17}\sqrt{u}*\frac{du}{32}\] \[\Large 2\pi*\frac{1}{32}\int_{1}^{17}\sqrt{u}du\] \[\Large \frac{\pi}{16}\int_{1}^{17}\sqrt{u}du\] \[\Large \frac{\pi}{16}*\frac{2}{3}u^{{}^{3/2}}+C {\Huge]}_{1}^{17}\] Plug in u = 17 and u = 1, then subtract. Do a bit of simplification to get \[\Large \frac{\pi(17\sqrt{17}-1)}{24}\]

jimthompson5910 (jim_thompson5910):

Notice how \(\Large \frac{\pi(17\sqrt{17}-1)}{24}\) is a single number and not a variable expression.

OpenStudy (clarence):

That's what I got if I were to integrate yours as well, right now I'm wondering why you got x where I got 4+2x^2

jimthompson5910 (jim_thompson5910):

I'm basing things off this page http://tutorial.math.lamar.edu/Classes/CalcII/SurfaceArea.aspx specifically the stuff in the blue box

OpenStudy (clarence):

My friends are being helpful this time and reckon you're right, so thanks for your help! Much appreciated!! :)

jimthompson5910 (jim_thompson5910):

you're welcome